OpenStudy (anonymous):

384 plus what gives you 96 does anyone know ill give a medal please help

OpenStudy (anonymous):

i meant 384 minus what gives you 96

OpenStudy (anonymous):

@wet

OpenStudy (anonymous):

384 - ? = 96 ? is some unknown number we'll be solving for.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Start by getting the variable to one of the sides.

OpenStudy (anonymous):

96

OpenStudy (anonymous):

384 - x = 96 Get x to one of the sides.

OpenStudy (anonymous):

the 96

OpenStudy (anonymous):

x is the variable we're solving for. Get x to one of the other sides. We're solving for x. Get x alone.

OpenStudy (anonymous):

then the 384

OpenStudy (anonymous):

use full sentences. I have no idea what you're trying to do

OpenStudy (anonymous):

whats the x for im confused

OpenStudy (anonymous):

x is the number we want to find. x is what we subtract from 384 to get 96

OpenStudy (anonymous):

oh ok i kinda get it

OpenStudy (anonymous):

So find x. 384 - x = 96

OpenStudy (anonymous):

i tryed using a calculator but the last number it pops up is 84

OpenStudy (anonymous):

384-x=96 -384 -384 -x = -288 x=288

OpenStudy (anonymous):

thank you so much :)

OpenStudy (anonymous):

don't just feed them answers...